# A multiple imputation approach to distinguish curative from life-prolonging effects in the presence of missing covariates

### v1

@@ 2.3 Multiple imputation for cure models
However, for the cure model, an additional challenge arises from the fact that even the cure status $G$ is unobserved for the censored observations. For a Cox PH cure model where the same design vector is common to both the incidence and the latency (i.e. $X=Z$), Beesley et al. (Beesley et al., 2016) derived the exact conditional distribution based on the complete likelihood function in (4). They proposed a chained equations procedure which imputes iteratively the missing values for the cure status $G$ and the incomplete covariates. In practice, drawing observations from the exact conditional distribution, which has a complicated form, is often computationally expensive. To reduce the computational time, Beesley et al. (Beesley et al., 2016) proposed approximations of the exact conditional distribution which are easier to handle in practice. However, assuming that $X$ and $Z$ are the same is quite restrictive, preventing us from fully exploiting the advantage of a mixture cure model in distinguishing between covariates with a curative or just a life-prolonging effect.
In this work, the ideas of Beesley et al. (Beesley et al., 2016) are extended to the more general model described in Section 2.1, considering three possible scenarios: the incomplete covariate belongs only to $X$, only to $Z$, or to both $X$ and $Z$. The scenario where the variable with missing values belongs to both $X$ and $Z$ is treated extensively in Section 3, while the cases where it belongs to only one of the two predictors (either the incidence or the latency component) are discussed in the appendices. Specifically, Appendix A deals with the case using the exact conditional distribution, while Appendix B addresses the case of the approximate conditional distribution. Further details about the method proposed by Beesley et al. (Beesley et al., 2016) and how it relates to this work are given in Section 3.
@@ 3 Methodology
In this section, we will describe our proposal for performing multiple imputation in cure models with incomplete covariates. As mentioned in Section 2.2, the main aspect of the multiple imputation procedure involves generating imputed values based on the conditional distribution of the incomplete variable given the other ones. Therefore, we will first derive the exact form of the conditional distribution and then outline the steps of the algorithm. Finally, we will introduce an approximation of the conditional distribution that enables us to generate imputed values at a reduced computational cost.
@@ 3.1 Using the exact conditional distribution
Let $W\in\mathbb{R}$ be the incomplete variable present in both the incidence and the latency model. The results for the cases where $W$ belongs either to the incidence or to the latency model are discussed in the Appendix. Since $W$ belongs to both $X$ and $Z$, there exist $1\leq j\leq p$ and $1\leq l\leq q$ such that $W=X^{(j)}=Z^{(l)}$. We denote the set of incidence and latency covariates excluding $W$ as $X^{(-j)}=X\setminus X^{(j)}$ and $Z^{(-l)}=Z\setminus Z^{(l)}$ respectively. To detail the proposed procedure in a more general context, we will consider two different choices for $W$, either a Gaussian or a Bernoulli random variable, conditional on the other covariates. However, results can be similarly generalized to other distributions.
@@
×$\quad\times\left[1-\pi(X_{i})\right]^{(1-\Delta_{i})(1-G_{i})}f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$
@@
×$\quad\times\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)$ (7)
Therefore, the exact distribution of $W$ given all the remaining covariates is as follows
$\begin{split}f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})&\propto\left(\frac{e^{\alpha_{0}+\alpha^{T}X_{i}}}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}e^{\beta^{T}Z_{i}\Delta_{i}}e^{-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}}\right)^{G_{i}}\\ &\quad\times\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)\end{split}$ (8)
One can draw values to impute $W$ from this distribution using the Metropolis-Hasting algorithm (Metropolis et al., 1953; Hastings, 1970). This algorithm can draw values from a given generic probability distribution (which we will call the target distribution) by sampling from a Markov chain whose stationary distribution is the target distribution.
Case 2. Alternatively, $W|X^{(-j)},Z^{(-l)}$ could be a categorical random variable. For instance, if $W$ follows a Bernoulli distribution, we can specify:
@@
where $\operatorname{expit}(x)=e^{x}/(1+e^{x})$ and $\mu$ is given by (6). The complete data likelihood is as in (3.1) where $f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ is replaced by $P(W_{i}=w_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ with $w_{i}\in\{0,1\}$. From it, and recalling that $\operatorname{logit}(x)=\log\bigl{(}\frac{x}{1-x}\bigr{)}$, the exact distribution of $W$ given all the remaining covariates takes the form
@@
$\qquad=G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\sum_{s\neq l}^{d}{\beta_{s}Z_{i,s}}}(e^{\beta_{l}}-1)+G_{i}\alpha_{j}+\log\left(1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)+$
$\qquad-\log\left(1+e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)+\mu_{i}$ (9)
Given that the conditional distributions depend on $\mu$ in both cases, one must draw values $\theta^{*}$ for the parameters $\theta$ before drawing values for the imputation of $W$. This is done by running a regression model (logistic regression when $W|X^{(-j)},Z^{(-l)}\sim$ Bernoulli and linear regression when $W|X^{(-j)},Z^{(-l)}\sim$ Normal), obtaining the point estimates $\hat{\theta}$ and the estimate for the covariance matrix $\hat{\Sigma}_{\theta}$ and drawing values $\theta^{*}$ from a Normal$\,(\hat{\theta},\hat{\Sigma}_{\theta})$ distribution.
@@ 3.2 Implementing multiple imputation in cure models
         a: Compute $\hat{\Psi}$ and $\hat{\Sigma}_{\Psi}$ from a logistic regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
         b: Draw values $\Psi^{*}$ from MVN$(\hat{\Psi},\hat{\Sigma}_{\Psi})$
         c: for all $W_{i}^{(0)}=$ NA do draw imputed values from $\operatorname{logit}[P(W_{i}=1|G_{i},\Delta_{i},Y_{i},X_{i}^{(-j)},Z_{i}^{(-l)};\Psi^{*})]$
@@
              a: Compute $\hat{\Psi}$ and $\hat{\Sigma}_{\Psi}$ from a linear regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
              b: Draw values $\Psi^{*}$ from MVN$(\hat{\Psi},\hat{\Sigma}_{\Psi})$
@@
3.3 Approximation using regression function
Drawing values from the exact conditional distributions in (8) and (3.1) can be very computationally intensive. Indeed, M-H algorithms are slowed down by the computation of complex target distributions and due to the slow convergence of such algorithms, one may need a very large number of iterations. Beesley et al. (Beesley et al., 2016) proposed an approximation and compared it to three alternatives from the literature. A simulation study indicated that the proposed approximation outperforms the others in terms of bias when the data has a cure fraction and $X=Z$. We extend this approach to cases where the sets of incidence and latency covariates differ. The following arguments will make use of the approximations
@@ 3.3 Approximation using regression function
First, if we assume that $W|X^{(-j)},Z^{(-l)}\sim N(\mu,\sigma)$, then the $\log$ of the exact conditional distribution of $W$ is given by
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $=G_{i}\Delta_{i}\beta_{l}W_{i}-G_{i}H_{0}(Y_{i})e^{\beta^{T}Z_{i}}+G_{i}\alpha_{j}W_{i}-\log\left(1+e^{\alpha_{0}+\alpha^{T}X_{i}}\right)+$
$-\frac{1}{2\sigma^{2}}\left(W_{i}-\mu_{i}\right)^{2}+C$
@@
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $\approx-\frac{1}{2\sigma^{2}}W_{i}^{2}+\left[G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\beta^{T}\bar{Z}}\beta_{l}+G_{i}\alpha_{j}+\right.$
$\quad\quad\left.-\frac{e^{\alpha_{0}+\alpha^{T}\bar{X}}}{1+e^{\alpha_{0}+\alpha^{T}\bar{X}}}\alpha_{j}+\frac{1}{\sigma^{2}}\mu_{i}\right]W_{i}+C$
Recalling that $\mu_{i}$ is a linear combination of $X_{i}^{(-j)}$ and $Z_{i}^{(-l)}$, one can see that the distribution of $W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i}$ is approximately Normal, where the mean is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$ and $G_{i}H_{0}(Y_{i})$. Therefore, a linear regression function with these covariates can be used to approximate the exact distribution of $W$.
@@
$\quad+G_{i}\alpha_{j}+\frac{e^{\alpha_{0}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}{1+e^{\alpha_{0}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}\left[1+\sum_{b\neq j}^{p}\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)\right]+$
$\quad-\frac{e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}{1+e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}\left[1+\sum_{b\neq j}^{p}\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)\right]+\mu_{i}+C$
This is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$ and $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$. Therefore, the exact distribution of $W$ can be approximated using a logistic regression function with these covariates as the last step of the imputation procedure presented in Section 3.2.
Details about the scenarios where the variable with missing values $W$ belongs only to the incidence or latency model are provided in Appendix B. If $W$ belongs only to the incidence, the conditional distribution of $W$ can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, and $G_{i}$. When $W$ belongs only to the latency, the conditional distribution of W can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$, and (for Case 2) $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$.

### v2

@@ 2.3 Multiple imputation for cure models
However, for the cure model, an additional challenge arises from the fact that even the cure status $G$ is unobserved for censored observations. For a Cox PH cure model where the design vector is common to both the incidence and the latency (i.e. $X=Z$), Beesley et al. (2016) derived the exact conditional distribution based on the complete likelihood function in (4). They proposed a chained equations procedure which imputes iteratively the missing values for the cure status $G$ and the incomplete covariates. In practice, drawing observations from the exact conditional distribution, which has a complicated form, is often computationally expensive. To reduce the computational time, Beesley et al. (2016) proposed approximations of the exact conditional distribution which are easier to handle in practice. However, assuming that $X$ and $Z$ are the same is quite restrictive, preventing us from fully exploiting the advantage of a mixture cure model in distinguishing between covariates with a curative or just a life-prolonging effect.
The ideas of Beesley et al. (2016) have been explicitly discussed in this limited context, but we think they can be extended to the more general model described in Section 2.1, considering three possible scenarios: the incomplete covariate belongs only to $X$, only to $Z$, or to both $X$ and $Z$. The scenario where the variable with missing values belongs to both $X$ and $Z$ is treated in Section 3, while the cases where it belongs to only one of the two components (either the incidence or the latency one) are discussed in the appendices. Specifically, Appendix A deals with the case using the exact conditional distribution, while Appendix B addresses the case of the approximate conditional distribution. Further details about the method proposed by Beesley et al. (2016) and how it relates to this work are given in Section 3.
@@ 3 Methodology
In this section, we describe the multiple imputation procedure in cure models with incomplete covariates. As mentioned in Section 2.2, the main aspect of the multiple imputation procedure involves generating imputed values based on the distribution of the incomplete variable conditional on the other ones. Therefore, we first present the exact form of the conditional distribution and then outline the steps of the algorithm. Finally, we introduce an approximation of the conditional distribution that enables us to generate imputed values at a reduced computational cost.
@@ 3.1 Using the exact conditional distribution
Let $W\in\mathbb{R}$ be the incomplete variable present in both the incidence and the latency model. The results for the cases where $W$ belongs either to the incidence or to the latency model are discussed in the Appendix. Since $W$ belongs to both $X$ and $Z$, there exist $1\leq j\leq p$ and $1\leq l\leq d$ such that $W=X^{(j)}=Z^{(l)}$. We denote the set of incidence and latency covariates excluding $W$ as $X^{(-j)}=X\setminus X^{(j)}$ and $Z^{(-l)}=Z\setminus Z^{(l)}$ respectively. To detail the proposed procedure in a more general context, we will consider two different choices for $W$, either having a Gaussian or a Bernoulli distribution, conditional on the other covariates. However, results can be similarly generalized to other distributions. We note also that the following formulas can be obtained from equation (2) in Beesley et al. (2016) by including all covariates to both parts of the model with corresponding coefficients equal to zero for the non-active covariates. However, in terms of estimation process, it makes a difference whether a certain covariate is included or not in each submodel.
@@
×$\,\times\left[1-\pi(X_{i})\right]^{(1-\Delta_{i})(1-G_{i})}f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$
@@
×$\,\times\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)$ (7)
Therefore, the exact distribution of $W$ conditional on all the remaining covariates is as follows
$f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\propto\left(\frac{e^{\alpha_{0}+\alpha^{T}X_{i}}}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}e^{\beta^{T}Z_{i}\Delta_{i}}e^{-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}}\right)^{G_{i}}\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)$ (8)
One can draw values to impute $W$ from this distribution using a Metropolis-Hasting algorithm (Metropolis et al., 1953; Hastings, 1970). This algorithm can draw values from a given generic probability distribution (the target distribution) by sampling from a Markov chain whose stationary distribution is the target distribution.
Case 2. Alternatively, $W|X^{(-j)},Z^{(-l)}$ could be a categorical random variable. For instance, if $W$ is binary, we can specify:
@@
where $\operatorname{expit}(x)=e^{x}/(1+e^{x})$ and $\mu$ is given by (6). The complete data likelihood is as in (3.1) where $f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ is replaced by $P(W_{i}=w_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ with $w_{i}\in\{0,1\}$. Recalling that $\operatorname{logit}(x)=\log\bigl{(}\frac{x}{1-x}\bigr{)}$, the exact distribution of $W$ conditional on all the remaining covariates takes the form
@@
$=G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\sum_{s\neq l}^{d}{\beta_{s}Z_{i,s}}}(e^{\beta_{l}}-1)+G_{i}\alpha_{j}$
$+\log\left(1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)-\log\left(1+e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)+\mu_{i}$ (9)
Given that the conditional distributions depend on $\mu$ in both cases, one must draw values $\theta^{*}$ for the parameters $\theta$ before drawing values for $W$. This is done by running a regression model (logistic regression when $W|X^{(-j)},Z^{(-l)}\sim$ Bernoulli and linear regression when $W|X^{(-j)},Z^{(-l)}\sim$ Normal), obtaining point estimates $\hat{\theta}$ and the corresponding covariance matrix estimates $\hat{\Sigma}_{\theta}$. Then, we draw values $\theta^{*}$ from a Normal$\,(\hat{\theta},\hat{\Sigma}_{\theta})$ distribution.
@@ 3.2 Implementing multiple imputation in cure models
         a: Compute $\hat{\theta}$ and $\hat{\Sigma}_{\theta}$ from a logistic regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
         b: Draw values $\theta^{*}$ from MVN$(\hat{\theta},\hat{\Sigma}_{\theta})$
         c: for all $W_{i}^{(0)}=$ NA do draw imputed values from $\operatorname{logit}[P(W_{i}=1|G_{i},\Delta_{i},Y_{i},X_{i}^{(-j)},Z_{i}^{(-l)};\theta^{*})]$
@@
              a: Compute $\hat{\theta}$ and $\hat{\Sigma}_{\theta}$ from a linear regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
              b: Draw values $\theta^{*}$ from MVN$(\hat{\theta},\hat{\Sigma}_{\theta})$
@@
3.3 Approximation of the conditional distribution
Drawing values from the exact conditional distributions in (8) and (3.1) can be computationally intensive. Indeed, M-H algorithms are slowed down by the computation of complex target distributions and due to the slow convergence of such algorithms, one may need a very large number of iterations. Beesley et al. (2016) proposed an approximation and compared it to three alternatives from the literature. A simulation study indicated that the proposed approximation outperforms the others in terms of bias when the data has a cure fraction and $X=Z$. We extend this approach to cases where the sets of incidence and latency covariates differ. As in Beesley et al. (2016), we make use of the first-order expansion
@@ 3.3 Approximation of the conditional distribution
First, we start discussing the case when $W|X^{(-j)},Z^{(-l)}\sim N(\mu,\sigma)$. The $\log$ of the exact conditional distribution of $W$ is given by
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $=G_{i}\Delta_{i}\beta_{l}W_{i}-G_{i}H_{0}(Y_{i})e^{\beta^{T}Z_{i}}+G_{i}\alpha_{j}W_{i}$
$-\log\left(1+e^{\alpha_{0}+\alpha^{T}X_{i}}\right)-\frac{1}{2\sigma^{2}}\left(W_{i}-\mu_{i}\right)^{2}+C$
@@
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $\approx-\frac{1}{2\sigma^{2}}W_{i}^{2}+\left[G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\beta^{T}\bar{Z}}\beta_{l}\right]W_{i}\tilde{C}$
$+\left[G_{i}\alpha_{j}-\frac{e^{\alpha_{0}+\alpha^{T}\bar{X}}}{1+e^{\alpha_{0}+\alpha^{T}\bar{X}}}\alpha_{j}+\frac{1}{\sigma^{2}}\mu_{i}\right]W_{i}\tilde{C}$
where $\tilde{C}$ is another constant with respect to $W_{i}$. Recalling that $\mu_{i}$ is a linear combination of $X_{i}^{(-j)}$ and $Z_{i}^{(-l)}$, one can see that the distribution of $W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i}$ is approximately Normal, where the mean is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$ and $G_{i}H_{0}(Y_{i})$. Then, instead of using the known expressions for the coefficients of this approximate linear approximation, we fit a linear regression model with these covariates to approximate the exact distribution of $W$.
@@
$+G_{i}\alpha_{j}+\frac{e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}{1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}\left[1+\sum_{b\neq j}^{p}{\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)}\right]$
$-\frac{e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}{1+e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}\left[1+\sum_{b\neq j}^{p}{\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)}\right]+\mu_{i}+C$
This is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$ and $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$. Therefore, the exact distribution of $W$ can be approximated by fitting a logistic regression model with these covariates as the last step of the imputation procedure presented in Section 3.2.
Details about the scenarios where the variable with missing values $W$ belongs only to the incidence or the latency model are provided in Appendix B. If $W$ belongs only to the incidence component, the conditional distribution of $W$ can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, and $G_{i}$. When $W$ belongs only to the latency, the conditional distribution of W can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$, and (for Case 2) $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$.
3.4 Variable selection in the presence of missing data
Variable selection in cure models has been addressed in the literature, with approaches such as the penalized likelihood methods proposed by Liu et al. (2012) and Masud et al. (2018). However, these methods do not account for missing data in covariates, leaving a critical gap in the current methodologies. Addressing this issue accurately is a complex challenge, which goes beyond the scope of this work. Here, we aim to provide preliminary suggestions by integrating a regularized Cox cure model (Masud et al., 2018) with guidelines from the literature on variable selection in the presence of missing data (Zhao and Long, 2017).
Statistical methods for variable selection combined with imputation generally fall into three primary strategies:
  • Impute-and-then-select: Apply a variable selection method (e.g., Lasso) to each imputed dataset separately and aggregate the selected variables across all imputed datasets.
  • Stacked imputation: Stack all imputed datasets into one combined dataset and perform variable selection as if the stacked data were a single dataset. This approach, while computationally efficient, may overstate the precision of variable selection by treating the stacked data as independent.
  • Bootstrap-and-impute: Combine bootstrap resampling with imputation to account for both the variability in the imputed datasets and the sampling variability in the variable selection process.
A commonly used method in the standard settings is the one proposed by Wood et al. (2008), which performs model selection on each imputed dataset separately. Hence, in the context of a mixture cure model, we propose to incorporate a regularized Cox cure model (Masud et al., 2018) into the multiple imputation process. Specifically, the variable selection process is repeated at each imputation step, using a penalized likelihood approach. The final set of selected variables is then defined as those selected in at least $\gamma M$ of the $M$ imputed datasets, where $\gamma$ is a user-defined threshold (e.g., $\gamma=0.5$).
Nonetheless, significant challenges remain. One major hurdle is the determination of tuning parameters, such as the penalization coefficient in Lasso, which typically requires cross-validation. Apart from being computationally intensive, cross-validation introduces additional variability and complicates the integration with multiple imputation. Moreover, the issue of post-selection inference - ensuring valid statistical inference after the variable selection process - remains unresolved in this context and is rarely addressed in existing studies on cure models. While this proposal provides a starting point for combining variable selection and imputation in cure models, a comprehensive investigation of these challenges is beyond the scope of this article. Our discussion here serves as an initial framework for future studies to build upon.